OpenStudy (anonymous):

A 2-column table with 'Type of Bag' and 'Fee per Bag' as the column headings. The first checked bag costs $20, the second checked bag costs $25, and extra bags cost $50 each If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all

OpenStudy (anonymous):

C = 145 20 + 25 +50(x-2) = 145 solve for x

OpenStudy (anonymous):

how do i solve for x

OpenStudy (anonymous):

20 + 25 +50(x-2) = 145 add like terms on left side 45 + 50(x-2) = 145 subtract 45 from both sides 50(x-2) = 100 divide both sides by 50 x - 2 = 2 add 2 to both sides x = 4

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

no prob